# The satisfiability threshold and solution space of random uniquely extendable constraint satisfaction problems

### v1

@@ 5 Non-group UE constraint functions
Suppose $k\geq 3$ and suppose $q=q_{1}q_{2}$, where $q_{1}\geq 3$. Suppose $q_{2}\geq 3$, or that $q_{2}=2$ and $k$ is odd. Then there is a $k$-ary commutative UE constraint on a spin set of size $q$ that is not equivalent to any group constraint.
@@
Let $q=q_{1}q_{2}$, where $q_{1}\geq 3$ and $q_{2}\geq 2$. Set $b=1$ if $k\neq 2\pmod{q_{2}}$ and set $b=0$ if $k=2\pmod{q_{2}}$. We consider the constraints $\psi_{\mathbb{Z}_{q_{1}},0}$ and $\psi_{\mathbb{Z}_{q_{2}},b}$, and let $\theta:\mathbb{Z}_{q_{1}}\to\Pi(\mathbb{Z}_{q_{2}})$ be such that $\theta_{(0,\dots,0)}$ is the transposition of 0 and $1$, and $\theta_{\sigma}=\operatorname{Id}$ for all $\sigma\in\mathbb{Z}_{q_{1}}^{k}\setminus\{(0,\dots,0)\}$. By Lemma 5.2, $\psi=\psi_{\mathbb{Z}_{q_{1}},0}\times_{\theta}\psi_{\mathbb{Z}_{q_{2}},1}$ is a $k$-ary commutative UE constraint on $\mathbb{Z}_{q_{1}}\times\mathbb{Z}_{q_{2}}$. We show that $\psi$ is not equivalent to any group constraint function. Define $f$ by
@@
Let $\eta\in(\mathbb{Z}_{q_{1}}\times\mathbb{Z}_{q_{2}})^{k-1}$ be such that $f_{\psi}(\eta)=(0,1)$. Let
$\sigma$ $=((-1,0),(1,0),(0,0),\ldots,(0,0))\in(\mathbb{Z}_{q_{1}}\times\mathbb{Z}_{q_{2}})^{k-1}$
$\sigma^{\prime}$ $=((1,0),(1,0),(0,0),\ldots,(0,0))\in(\mathbb{Z}_{q_{1}}\times\mathbb{Z}_{q_{2}})^{k-1}$
$\sigma^{\prime\prime}$ $=((-1,0),(-1,0),(0,0),\ldots,(0,0))\in(\mathbb{Z}_{q_{1}}\times\mathbb{Z}_{q_{2}})^{k-1}.$
@@
If $b=1$ then $k\neq 2\pmod{q_{2}}$, so $b+3-k=2-k\neq 0\pmod{q_{2}}$. If $b=0$ then $k=2\pmod{q_{2}}$ and moreover $q_{2}\geq 3$, as the case $q_{2}=2$ and $k$ being even is excluded by the hypotheses of the theorem. In this case we have $b+3-k=3-k=1\pmod{q_{2}}$. If $q_{2}\geq 4$ then $\theta_{(0,\ldots,0)}^{-1}(-2)=-2\neq 1\pmod{q_{2}}$. If $q_{2}=3$, then $\theta_{(0,\ldots,0)}^{-1}(-2)=0\neq 1\pmod{q_{2}}$. Therefore $f(\sigma^{\prime}\sigma^{\prime\prime}\eta^{k-3})\neq f(\sigma\sigma\eta^{k-3})$. ∎
@@ Acknowledgement
Here we briefly sketch how a special case of Theorem 1.16 when $\Psi=\{\psi\}$, restated below, can be alternatively proved using quasigroup theory. Recall that $f^{(i)}_{\psi}$ denotes the unique function in $F_{i}(\{\psi\})$.
@@
Let $\psi$ be a $k$-ary UE commutative constraint function with $k\geq 3$. Then $\psi$ is a group constraint if and only if $f^{(2)}_{\psi}$ is a symmetric function.
A Symmetry
@@
B Reducibility
We sketch an alternative proof for the special case of Theorem 1.17 when $\Psi=\{\psi\}$ using “reducibility” of quasigroups. In quasigroup theory, the term reducibility refers to a notion that differs from the one we introduced in Definition 1.9. To avoid confusion, we will refrain from using the word reducibility by itself. Instead, we introduce the terms exchangeability and group-reducibility: Exchangeability corresponds exactly to the notion defined as reducibility in Definition 1.9, whereas group-reducibility refers to the standard notion of reducibility used in quasigroup theory.
@@ B Reducibility
For every completely group-reducible totally symmetric (k1)(k-1)-quasigroup $(\Omega,f)$ with $k-1\geq 3$, there is an Abelian group $(Q,+)$, a permutation $\varphi$ of $\Omega$, and $b\in\Omega$ such that, for all $\sigma=\sigma_{1}\ldots\sigma_{k-1}\in\Omega^{k-1}$,
@@
Proof of the special case of Theorem 1.17 for $\Psi=\{\psi\}$. First suppose that there is a group $(\Omega,+)$ and $b\in\Omega$ such that $\psi=\psi_{(\Omega,+),b}$. To show that $\{\psi\}$ is reducible, let $d\in\{0,\ldots,k\}$ and let $\sigma,\sigma\in\Omega^{d}$, $\eta,\eta^{\prime}\in\Omega^{k-d}$ be such that
@@
which implies $\sigma^{\oplus}=\sigma^{\prime\oplus}$ and $\eta^{\oplus}=\eta^{\prime\oplus}$. Thus $\sigma^{\prime\oplus}+\eta^{\prime\oplus}=\sigma^{\oplus}+\eta^{\oplus}=b$, and therefore $\sigma^{\prime}\eta^{\prime}\models\psi$. Therefore $\{\psi\}$ is reducible.
For the converse direction, suppose $\{\psi\}$ is reducible. By Theorem B.5, it suffices to show that the (k1)(k-1)-quasigroup $(\Omega,f_{\psi})$ is totally symmetric and completely group-reducible. To show that $(\Omega,f_{\psi})$ is totally symmetric, let $\sigma_{1},\ldots,\sigma_{k}\in\Omega^{k}$, and let $\pi\in\Pi([k])$. Suppose $f_{\psi}(\sigma_{1},\ldots,\sigma_{k_{1}})=\sigma_{k}$. Then we have
@@
Therefore, by reducibility of $\{\psi\}$, we can replace $\eta$ in $\eta\sigma_{1}^{\prime}\sigma_{2}^{\prime}f_{\psi}(\sigma_{1}^{\prime}\eta\sigma_{2}^{\prime})\models\psi$ with $\eta^{\prime}$ to get $\eta^{\prime}\sigma_{1}^{\prime}\sigma_{2}^{\prime}f_{\psi}(\sigma_{1}^{\prime}\eta\sigma_{2}^{\prime})\models\psi$. Applying commutativity again gives

### v2

@@ 5 Non-group UE constraint functions
Suppose $k\geq 3$ and suppose $r=r_{1}r_{2}$, where $r_{1}\geq 3$. Suppose $r_{2}\geq 3$, or that $r_{2}=2$ and $k$ is odd. Then there is a $k$-ary commutative UE constraint on a spin set of size $r$ that is not equivalent to any group constraint.
@@
Let $r=r_{1}r_{2}$, where $r_{1}\geq 3$ and $r_{2}\geq 2$. Set $b=1$ if $k\neq 2\pmod{r_{2}}$ and set $b=0$ if $k=2\pmod{r_{2}}$. We consider the constraints $\psi_{\mathbb{Z}_{r_{1}},0}$ and $\psi_{\mathbb{Z}_{r_{2}},b}$, and let $\theta:\mathbb{Z}_{r_{1}}\to\Pi(\mathbb{Z}_{r_{2}})$ be such that $\theta_{(0,\dots,0)}$ is the transposition of 0 and $1$, and $\theta_{\sigma}=\operatorname{Id}$ for all $\sigma\in\mathbb{Z}_{r_{1}}^{k}\setminus\{(0,\dots,0)\}$. By Lemma 5.2, $\psi=\psi_{\mathbb{Z}_{r_{1}},0}\times_{\theta}\psi_{\mathbb{Z}_{r_{2}},1}$ is a $k$-ary commutative UE constraint on $\mathbb{Z}_{r_{1}}\times\mathbb{Z}_{r_{2}}$. We show that $\psi$ is not equivalent to any group constraint function. Define $f$ by
@@
Let $\eta\in(\mathbb{Z}_{r_{1}}\times\mathbb{Z}_{r_{2}})^{k-1}$ be such that $f_{\psi}(\eta)=(0,1)$. Let
$\sigma$ $=((-1,0),(1,0),(0,0),\ldots,(0,0))\in(\mathbb{Z}_{r_{1}}\times\mathbb{Z}_{r_{2}})^{k-1}$
$\sigma^{\prime}$ $=((1,0),(1,0),(0,0),\ldots,(0,0))\in(\mathbb{Z}_{r_{1}}\times\mathbb{Z}_{r_{2}})^{k-1}$
$\sigma^{\prime\prime}$ $=((-1,0),(-1,0),(0,0),\ldots,(0,0))\in(\mathbb{Z}_{r_{1}}\times\mathbb{Z}_{r_{2}})^{k-1}.$
@@
If $b=1$ then $k\neq 2\pmod{r_{2}}$, so $b+3-k=2-k\neq 0\pmod{r_{2}}$. If $b=0$ then $k=2\pmod{r_{2}}$ and moreover $r_{2}\geq 3$, as the case $r_{2}=2$ and $k$ being even is excluded by the hypotheses of the theorem. In this case we have $b+3-k=3-k=1\pmod{r_{2}}$. If $r_{2}\geq 4$ then $\theta_{(0,\ldots,0)}^{-1}(-2)=-2\neq 1\pmod{r_{2}}$. If $r_{2}=3$, then $\theta_{(0,\ldots,0)}^{-1}(-2)=0\neq 1\pmod{r_{2}}$. Therefore $f(\sigma^{\prime}\sigma^{\prime\prime}\eta^{k-3})\neq f(\sigma\sigma\eta^{k-3})$. ∎
@@ Acknowledgement
Here we briefly sketch how the special cases of Theorems 1.16 and 1.17 when $\Psi=\{\psi\}$, restated below, can be alternatively proved using quasigroup theory. Recall that $f^{(i)}_{\psi}$ denotes the unique function in $F_{i}(\{\psi\})$.
@@
Let $\psi$ be a $k$-ary UE commutative constraint function with $k\geq 3$. Then $\psi$ is equivalent to a group constraint function if and only if $f^{(2)}_{\psi}$ is a symmetric function.
Theorem 5.5.
Let $\psi$ be a $k$-ary UE commutative constraint function with $k\geq 4$. Then $\psi$ is equivalent to a group constraint function if and only if $\psi$ is reducible.
A Symmetry and proof of Theorem 5.4
@@
B Reducibility and proof of Theorem 5.5
We sketch an alternative proof for the special case of Theorem 1.17 when $\Psi=\{\psi\}$ using “reducibility” of quasigroups. In quasigroup theory, the term reducibility refers to a notion that differs from the one we introduced in Definition 1.9. To avoid confusion, we will refrain from using the word reducibility by itself when we discuss quasigroups. Instead, we introduce the terms exchangeability and group-reducibility: Exchangeability for quasigroups, defined below, corresponds exactly to the notion defined as reducibility in Definition 1.9 for UE constraint functions, whereas group-reducibility refers to the standard notion of reducibility in quasigroup theory.
@@
For every completely group-reducible totally symmetric (k1)(k-1)-quasigroup $(\Omega,f)$ with $k-1\geq 3$, there is an Abelian group $(\Omega,+)$, a permutation $\varphi$ of $\Omega$, and $b\in\Omega$ such that, for all $\sigma=\sigma_{1}\ldots\sigma_{k-1}\in\Omega^{k-1}$,
@@
Proof of Theorem 5.5. First suppose that there is a group $(\Omega,+)$ and $b\in\Omega$ such that $\psi=\psi_{(\Omega,+),b}$. To show that $\psi$ is reducible, let $d\in\{0,\ldots,k\}$ and let $\sigma,\sigma\in\Omega^{d}$, $\eta,\eta^{\prime}\in\Omega^{k-d}$ be such that
@@
which implies $\sigma^{\oplus}=\sigma^{\prime\oplus}$ and $\eta^{\oplus}=\eta^{\prime\oplus}$. Thus $\sigma^{\prime\oplus}+\eta^{\prime\oplus}=\sigma^{\oplus}+\eta^{\oplus}=b$, and therefore $\sigma^{\prime}\eta^{\prime}\models\psi$. Therefore $\psi$ is reducible.
For the converse direction, suppose $\{\psi\}$ is reducible. By Theorem B.5, it suffices to show that the (k1)(k-1)-quasigroup $(\Omega,f_{\psi})$ is totally symmetric and completely group-reducible. To show that $(\Omega,f_{\psi})$ is totally symmetric, let $\sigma_{1},\ldots,\sigma_{k}\in\Omega^{k}$, and let $\pi\in\Pi([k])$. Suppose $f_{\psi}(\sigma_{1},\ldots,\sigma_{k-1})=\sigma_{k}$. Then we have
@@
Therefore, by reducibility of $\{\psi\}$ and Remark 1.10, we can replace $\eta$ in $\eta\sigma_{1}^{\prime}\sigma_{2}^{\prime}f_{\psi}(\sigma_{1}^{\prime}\eta\sigma_{2}^{\prime})\models\psi$ with $\eta^{\prime}$ to get $\eta^{\prime}\sigma_{1}^{\prime}\sigma_{2}^{\prime}f_{\psi}(\sigma_{1}^{\prime}\eta\sigma_{2}^{\prime})\models\psi$. Applying commutativity again gives